# Noemi Walczak Université Paris 1 Panthéon-Sorbonne, Paris, France Kenan Huremović Armando Rungi

### v3

@@ 2. INSTITUTIONAL SETTING
Global greenhouse gas (GHG) emissions have increased steadily since the beginning of the 21st century, despite international efforts to curb them [14]. In response, the European Union has committed to reducing the net domestic greenhouse gas emissions by at least 55% by 2030 relative to 1990 levels, and achieving climate neutrality by 2050 [16]. Central to this strategy is the EU Emission Trading Scheme (ETS), [^2] The ETS is also in place in Norway, Liechtenstein, Northern Ireland, and Iceland, the first major carbon pricing market. The ETS is a cap-and-trade mechanism designed to reduce GHG emissions. It sets a limit (cap) on the total amount of emissions that can be produced by facilities covered under the system. To help achieve the EU’s climate goals, this cap decreases annually, ensuring a gradual and consistent reduction in overall emissions. Within the emissions cap, allowances, measured in tons of $\text{CO}_{2}$ equivalent, are distributed through a combination of auctioning and free allocation. The system covers around 40% of European GHG emissions and, since its launch in 2005, it has helped drive down emissions from electricity, heat generation and industrial production by 47% [17] [^3] See Appendix Figure A.3 for more details..
@@
One consequence of the ETS is that producers covered by the system may face a competitive disadvantage compared to similar producers in countries without a comparable carbon pricing scheme. This creates an incentive for EU firms to relocate carbon-intensive stages of production to countries with less stringent environmental regulations—a phenomenon commonly referred to as carbon leakage. As an illustration, Figure 1 shows the positive correlation between each non-ETS country’s weighted average carbon intensity and their integration into the European supply network. We compute the weighted average carbon intensity by assigning to each sectoral pollution intensity a weight corresponding to the sector’s share of total country’s sales. Integration into the European supply network corresponds, instead, to the share of both direct and indirect exports to the EU, relative to the country’s total sales.
Figure 1: Carbon Intensity and Trade Integration into the EU supply network of non-ETS trading partners
@@
The positive correlation underlines an important feature of the EU’s trade pattern: it is highly integrated with countries that tend to have emission-intensive production structures. This pattern creates a misalignment between EU’s climate goals and the carbon content of its imports, undermining European’s internal efforts by ”offshoring” emissions through the purchase of carbon-intensive goods from foreign countries.
@@
To address these concerns, in 2021, the European Commission proposed the introduction of the Carbon Border Adjustment Mechanism (CBAM), complementing it with the gradual phase-out of the free allowances. The CBAM is a policy tool designed to place a carbon price on the emissions embodied in imports of specific goods from non-EU countries, ensuring that importers face similar carbon costs as those borne by EU producers sourcing the same goods domestically. This aims to prevent carbon leakage and maintain the competitiveness of EU industries under the ETS. Under the CBAM, European importers are required to purchase and submit CBAM certificates corresponding to the embedded emissions in their imported goods. The price of these certificates is based on the weekly average closing price of EU ETS allowances on the auction platform. This mechanism ensures that EU producers continue to face carbon costs under the EU ETS, while foreign exporters of emissions-intensive goods are subject to a comparable carbon price, thereby leveling the playing field and reducing the risk of carbon leakage. By introducing a carbon price on the carbon content of imports, CBAM serves as a hybrid instrument of trade and climate policy, safeguarding the EU’s climate targets from being undermined by emissions embodied in imports.
@@
Since 2023, the CBAM has been in a transitional phase, requiring firms in selected sectors to report the embodied emissions of their imports without financial obligations. Starting in 2026, importers will be required to purchase CBAM certificates corresponding to the carbon content of covered goods (iron and steel, cement, aluminum, organic basic chemicals, hydrogen, fertilizers, and electricity). The mechanism is expected to expand to all imported goods corresponding to those falling under the ETS scheme by 2030.
In the following section, we develop a model to assess the impact of the CBAM on trade and emissions embodied in trade flows, and quantify its effectiveness as both a climate and competitiveness policy instrument.
@@ 3. THE MODEL
We develop a multi-country, multi-sector model with country-sectoral linkages and trade in intermediate goods, building on [7]. There is a set $\mathcal{N}=\{1,...,N\}$ of countries, indexed by $i$ and $n$, and a set $\mathcal{J}=\{1,...,J\}$ of sectors for each country, indexed by $j$ and $k$. We use subscripts to denote countries, and superscripts to denote sectors.
@@
Each sector $j$ in country $i$ is represented by a competitive producer that produces a single good using labor $l_{i}^{j}$ and materials $M_{i}^{j}$ as inputs. Each sectoral good can be either consumed as a final good by the household or used as an intermediate input in the production of other goods. Since each sector produces one good, we will use terms good/input and sector interchangeably when there is no fear of ambiguity. Following [11], we assume that a fraction $a_{i}^{j}\in(0,1)$ of input mix is used to abate pollution. We use $q_{i}^{j}$ to denote the output of sector $j$ in country $i$, and assume that producers combine inputs in a Cobb-Douglas fashion. The production function exhibits constant returns to scale and has the following form:
@@
where $A_{i}^{j}$ is a Hicks-neutral parameter, $l_{i}^{j}$ denotes the labor input, and $\beta_{i}^{j}>0$ is its share in production. $\Upsilon_{i}^{j}$ is a normalizing constant, that we specify later. Moreover, we define materials $M_{i}^{j}$ as a CES aggregate of intermediate goods used in the production, namely:
@@
We assume that each sector generates emissions $e_{i}^{j}$ as a by-product. The amount of emissions negatively depends on the abatement, with $\rho_{i}^{j}$ being the emissions elasticity:
@@
Given the functional form of the emissions’ function, under the assumption that $a_{i}^{j}\in(0,1)$, there is a one-to-one mapping between the emissions $e_{i}^{j}$ and the abatement $a_{i}^{j}$. In the following, we treat $e_{i}^{j}$ as the producers’ choice variable and use equation (4) as our production function. The same approach is taken, for instance, in [27].
@@
Following [3], we distinguish between countries that have a national carbon market (belonging to the set $\mathcal{N}_{c}$) and countries that have not adopted any pricing mechanism on emissions ($\mathcal{N}_{nc}$). For the first group, we consider a national inelastic supply of emissions, and the unit price of emissions $t_{i}$ is determined endogenously to equate the aggregate demand for emissions with the available supply. This price can be interpreted as the shadow price of all regulations applying a price on emissions. Specifically, in cap-and-trade systems, total national emissions are limited by the total number of allowances, each of which permits the emission of one unit $e_{i}^{j}$ of polluting emissions. Accordingly, $t_{i}$ represents the market price of these emission permits, which may either be purchased or allocated for free. We let $\epsilon_{i}^{j}\in(0,1)$ denote the share of emission permits that are freely distributed to a sector $j$ in country $i$, relative to the total number of permits, and adjust national supply by discounting for the free allowances. In contrast, for countries in $\mathcal{N}_{nc}$, we impose an exogenous carbon price. In this case, the supply of emissions permits is endogenously determined to match the demand at the given price.
@@
We assume that trade across countries and sectors is costly and subject to frictions modeled as a combination of iceberg trade costs $d_{ni}^{kj}\geq 1$, with $d_{ii}^{kj}=1\ \forall i,n,k,j$, and ad-valorem tariffs $\kappa_{ni}^{kj}$ [^4] We retain the destination sector index $j$ to capture potential heterogeneity in tariffs across importing sectors and to maintain flexibility for more realistic modeling.. Importantly, we directly incorporate the Carbon Border Adjustment Mechanism (CBAM) that targets the emissions embodied in trade, as an additional trade friction. Recall that the CBAM aims to ensure that imported goods face a comparable carbon cost to domestically produced goods, to mitigate carbon leakage and level the playing field.
@@
To simplify notation, we define $\tilde{\tau}_{ni}^{kj}\equiv 1+\kappa_{ni}^{kj}+CBAM_{ni}^{kj}$. Our modeling approach is coherent with the predominant approach in literature that models carbon border adjustments as an additional tariff [^5] Recent examples modeling the European CBAM as an additional tariff include [3, 8, 23, 19, 12]..
We examine two distinct approaches to determining the value of CBAM, which we label exogenous and endogenous CBAM. When the value of CBAM is set based on the prevailing vector of emission prices $\mathbf{t}$, we refer to this as the exogenous case. However, because the introduction of CBAM generally alters the relative carbon prices across countries, we define the endogenous CBAM as the value that emerges when emission prices $\mathbf{t}$ adjust in response to the implementation of CBAM. This is a distinguishing feature of our approach and is not incorporated in related work.
@@
$c_{i}^{j}=\frac{I_{i}}{\sum_{k\in\mathcal{J}}(\chi_{i}^{k}/\chi_{i}^{j})(p_{i}^{k})^{1-\sigma}(p_{i}^{j})^{\sigma}}\quad\forall i\in\mathcal{N},j\in\mathcal{J}$
@@
We begin by defining the equilibrium in the case where the CBAM value is fixed based on the initial vector of emission prices $\mathbf{t}$ and does not respond to subsequent price changes triggered by the introduction of CBAM, as discussed earlier in this Section. We then extend this definition to allow the CBAM value to adjust endogenously in response to the price shifts induced by the trade barrier it introduces.
@@
Our model conceptualizes the world economy as a network, where each node represents a sector in a country, and the links denote the flows of intermediate inputs between these sector-country pairs. One of the goals of our analysis is to evaluate the role of the structure of this network in mediating the effects of the CBAM. To ease the analysis of the model, we introduce additional notation.
To facilitate calibration, we define the steady state of the economy as the situation in which wedges $d_{ni}^{kj}$ and $\kappa_{ni}^{kj}$ are equal to zero for all combinations of $i$, $n$, $k$, $j$. Furthermore, we normalize the hicks productivity parameter $A_{i}^{j}$ such that ${(1-\epsilon_{i}^{j})}^{\rho_{i}^{j}}{A_{i}^{j}}^{-(1-\rho_{i}^{j})}=1.$ [^6] Note that this is equivalent to appropriately rewriting the normalization constant $\Upsilon_{i}^{j}$ in (5).
@@
    $\alpha_{i}^{j}\equiv\frac{p_{i}^{j}c_{i}^{j}}{\sum_{j\in\mathcal{J}}p_{i}^{j}c_{i}^{j}}=\frac{p_{i}^{j}c_{i}^{j}}{I_{i}}=\chi_{i}^{j}\frac{(p_{i}^{j})^{1-\sigma}}{\sum_{j\in\mathcal{J}}\chi_{i}^{j}(p_{i}^{j})^{1-\sigma}}$
@@
4. QUANTIFYING THE EFFECT OF THE CBAM
In this section, we apply the model from the previous section to evaluate the effects of the introduction of the CBAM on trade across sectors and countries, emissions embodied in imports, carbon leakage, gross national expenditure, and real wages.
We solve the model in changes using hat algebra approach ([13]), enabling us to evaluate different counterfactual scenarios. A desirable feature of this approach is that it captures the full non-linear adjustments to a finite – though not necessarily marginal – policy shock, while keeping computational efficiency. To complement this analysis and to illustrate the mechanisms at play, we provide comparative statics analysis of the effects of a marginal adjustment in CBAM in the Appendix A.2.
In our analysis, we consider both the adjustment to the equilibrium with exogenous CBAM and endogenous CBAM, as explained in the previous section. Unlike standard trade tariffs, which are fixed, the CBAM adjusts in response to domestic and foreign carbon prices – both of which may be influenced by the implementation of CBAM itself. By considering both cases, we can assess the significance of this additional adjustment mechanism. Furthermore, we distinguish between the current CBAM and a full CBAM. As noted in the introduction, the current CBAM applies only to a subset of goods, which in our analysis correspond to six sectors: Mining and Quarrying; Chemicals and Chemical Products; Non-Metallic Mineral Products; Basic Metals; Fabricated Metal Products; and Electricity, Gas, Steam, and Air Conditioning Supply. In contrast, the full CBAM extends coverage to all sectors of the EU economy currently covered by the ETS.

### v4

@@ 3. Institutional setting
Global greenhouse gas (GHG) emissions have increased steadily since the beginning of the 21st century, despite international efforts to curb them [ipccreport]. In response, the European Union has committed to reducing the net domestic greenhouse gas emissions by at least 55% by 2030 relative to 1990 levels, and achieving climate neutrality by 2050 [FitFor55regulation]. Central to this strategy is the EU Emission Trading Scheme (ETS), the first major carbon pricing market, which also applies to Norway, Liechtenstein, Northern Ireland, and Iceland. The ETS is a cap-and-trade mechanism: it sets a limit (cap) on a decreasing total amount of emissions that can be produced to achieve EU climate goals, then a corresponding amount of allowances, measured in tons of $\text{CO}_{2}$ equivalent, are distributed through a combination of auctioning and free allocation [^2] Free allocation mainly targets sectors and sub-sectors that are considered most at risk of carbon leakage, in order to prevent the relocation of production to regions with less stringent climate policies. After the actual transitional phase, the share of free allowances should decrease while CBAM certificates become available..
The system covers around 40% of European GHG emissions and, since its launch in 2005, it has helped drive down emissions from electricity, heat generation, and industrial production by 47% [EUreport2023] [^3] See Appendix Figure A.3 for more details..
@@
One consequence of the ETS is that producers covered by the system may face a competitive disadvantage compared to similar producers in countries without a comparable carbon pricing scheme. This creates an incentive for EU firms to relocate carbon-intensive stages of production to countries with less stringent environmental regulations—a phenomenon commonly referred to as carbon leakage. As an illustration, Figure 1 shows the positive correlation between each non-ETS country’s weighted average carbon intensity and their integration into the European supply network. We calculate the weighted average carbon intensity by assigning a weight to each sectoral carbon intensity that corresponds to the sector’s share of the total country’s sales. Integration into the European supply network corresponds, instead, to the share of both direct and indirect exports to the EU, relative to the country’s total sales.
Figure 1: Carbon Intensity and Trade Integration into the EU supply network of non-ETS trading partners. Carbon Intensity is computed as the weighted average carbon intensity of each country, where the weights corresponds to the sector’s share of the total country’s sales. Trade Integration, instead, is computed as the country share of both direct and indirect exports to the EU, relative to the country’s total sales.
@@
The positive correlation underlines an important feature of the EU trade pattern: it is relatively highly integrated with countries that tend to have emission-intensive production structures. This pattern creates a misalignment between EU climate goals and the carbon content of its imports, undermining Europe’s internal efforts by ”offshoring” emissions through the purchase of carbon-intensive goods from foreign countries [^4] After a second observation, Figure 1 primarily underscores the heterogeneity in carbon intensity across EU partners. For more details, see also Appendix Table A2.
@@
To address these concerns, in 2021, the European Commission proposed the introduction of the Carbon Border Adjustment Mechanism (CBAM), complementing it with the gradual phase-out of the free allowances (Regulation (EU) 2023/956), which were previously provided to prevent moving production to countries with less strict carbon regulations. The CBAM is a policy tool designed to place a carbon price on the emissions embodied in imports of specific goods from non-EU countries, ensuring that importers face similar carbon costs as those borne by EU producers sourcing the same goods domestically. This aims to prevent carbon leakage and maintain the competitiveness of EU industries under the ETS. Under the CBAM, European importers are required to purchase and submit CBAM certificates corresponding to the embedded emissions in their imported goods. The price of these certificates is based on the weekly average closing price of EU ETS allowances on the auction platform. This mechanism ensures that EU producers continue to face carbon costs under the EU ETS, while foreign exporters of emissions-intensive goods are virtually subject to a comparable carbon price, thereby leveling the playing field and reducing the risk of carbon leakage. By introducing a carbon price on the carbon content of imports, CBAM serves as a hybrid instrument of trade and climate policy, safeguarding the EU climate targets from being undermined by emissions embodied in imports.
Yet, two major downsides emerge. The first is that the design of a border adjustment mechanism, such as CBAM, discriminates across countries because the price of the same input varies depending on the origin country’s level of emissions for that product. This is not compatible with the WTO clause on the Most Favored Nation (MFN), which implies that the same good imported from any two different origin countries must have the same tariffs or non-tariff barriers as in the most favored trade partner country. The point of view by the European Commission is that, once one evaluates the amount of regulatory burden, there is no discrimination between EU producers and any other partner country, and the mechanism is intended to equalize an otherwise unequal regulatory burden, therefore respecting both WTO principles of National Treatment (NT) and Most Favored Nation (MFN).
@@
The second downside is that a well-functioning CBAM requires the establishment of an information system to collect and update data on emissions for different products in different countries. It may not be trivial to establish such a system in time for the end of the transitional phase. In fact, since 2023, the CBAM has been in a transitional phase, requiring firms in selected sectors to report the embodied emissions of their imports without financial obligations. Starting in 2026, importers will be required to purchase CBAM certificates corresponding to the carbon content of covered goods (iron and steel, cement, aluminum, organic basic chemicals, hydrogen, fertilizers, and electricity). The mechanism is expected to expand to all imported goods corresponding to those falling under the ETS scheme by 2030. By that time, a proper informative system is required to work at full capacity.
@@ 4. The model
We develop a multi-country, multi-sector model with country-sectoral linkages, building on [caliendo2022distortions]. There is a set $\mathcal{N}=\{1,...,N\}$ of countries, indexed by $i$ and $n$, and a set $\mathcal{J}=\{1,...,J\}$ of sectors for each country, indexed by $j$ and $k$. We use subscripts to denote countries, and superscripts to denote sectors.
@@
Each sector $j$ in country $i$ is represented by a competitive producer that produces a single good using labor $l_{i}^{j}$ and materials $M_{i}^{j}$ as inputs. Each sectoral good can be either consumed as a final good by the household or used as an intermediate input in the production of other goods. Since we assume that each sector produces one good, we will use terms good/input and sector interchangeably when there is no fear of ambiguity. Following [copeland2003trade], we assume that a fraction $a_{i}^{j}\in(0,1)$ of the input mix is used to abate pollution. We use $q_{i}^{j}$ to denote the output of sector $j$ in country $i$, and assume that producers combine inputs in a Cobb-Douglas fashion. The production function exhibits constant returns to scale and has the following form:
@@
where $A_{i}^{j}$ is a Hicks-neutral productivity parameter, $l_{i}^{j}$ denotes the labor input, and $\beta_{i}^{j}>0$ is its share in production. $\Upsilon_{i}^{j}$ is a normalizing constant that we specify later. Moreover, we define materials $M_{i}^{j}$ as a CES aggregate of intermediate goods used in the production, namely:
@@
Crucially, we assume that each sector generates emissions $e_{i}^{j}$ as a by-product. The amount of emissions negatively depends on the abatement, with $\rho_{i}^{j}$ being the emissions elasticity:
@@
Given the functional form of the emissions’ function, under the assumption that $a_{i}^{j}\in(0,1)$, there is a one-to-one mapping between the emissions $e_{i}^{j}$ and the abatement $a_{i}^{j}$. In the following, we treat $e_{i}^{j}$ as the producers’ choice variable and use equation (4) as our production function. The same approach is taken, for instance, in [shapiro2018pollution].
@@
Following [bellora2023eu], we distinguish between countries that have a national carbon market (belonging to the set $\mathcal{N}_{c}$) and countries that have not adopted any pricing mechanism on emissions ($\mathcal{N}_{nc}$). For the first group, we consider a national inelastic supply of emissions, and the unit price of emissions $t_{i}$ is determined endogenously to equate the aggregate demand for emissions with the available supply. This price can be interpreted as the shadow price of all regulations applying a price on emissions. Specifically, in cap-and-trade systems, total national emissions are limited by the total number of allowances, each of which permits the emission of one unit $e_{i}^{j}$ of polluting emissions. Accordingly, $t_{i}$ represents the market price of these emission permits, which may either be purchased or allocated for free. We let $\epsilon_{i}^{j}\in(0,1)$ denote the share of emission permits that are freely distributed to a sector $j$ in country $i$, relative to the total number of permits, and adjust national supply by discounting for the free allowances. In contrast, for countries in $\mathcal{N}_{nc}$, we impose an exogenous carbon price. In this case, the supply of emissions permits is endogenously determined to match the demand at the given price.
@@
We assume that trade across countries and sectors is costly and subject to frictions modeled as a combination of iceberg trade costs $d_{ni}^{kj}\geq 1$, with $d_{ii}^{kj}=1\ \forall i,n,k,j$, and ad-valorem tariffs $\kappa_{ni}^{kj}$ [^5] We retain the destination sector index $j$ to capture potential heterogeneity in tariffs across importing sectors and to maintain flexibility for more realistic modeling. A similar approach is adopted, for example, in [baqaee2024networks].. Importantly, we directly incorporate the Carbon Border Adjustment Mechanism (CBAM) that targets the emissions embodied in trade, as an additional trade friction. Recall that the CBAM aims to ensure that imported goods face a comparable carbon cost to domestically produced goods, to mitigate carbon leakage and level the playing field.
@@
To simplify notation, we define $\tilde{\tau}_{ni}^{kj}\equiv 1+\kappa_{ni}^{kj}+CBAM_{ni}^{kj}$. Our modeling approach is coherent with the predominant approach in literature that models carbon border adjustments as an additional tariff [^6] Recent examples modeling the European CBAM as an additional tariff include [bellora2023eu, campolmi2024designing, larch2024consequences, florez2024eu, coster2024firms]..
We examine two distinct approaches to determining the value of CBAM, which we label exogenous and endogenous CBAM, respectively. When the value of CBAM is set based on the prevailing vector of emission prices $\mathbf{t}\in\mathbb{R}^{N}$, we refer to this as the exogenous case. However, because the introduction of CBAM generally alters the relative carbon prices across countries, we define the endogenous CBAM as the value that emerges when emission prices $\mathbf{t}$ adjust in response to the implementation of CBAM. The latter is a distinguishing feature of our approach that we haven’t found in related work.
@@
$c_{i}^{j}=\frac{I_{i}}{\sum_{k\in\mathcal{J}}(\chi_{i}^{k}/\chi_{i}^{j})(p_{i}^{k})^{1-\sigma}(p_{i}^{j})^{\sigma}}\qquad\forall i\in\mathcal{N},j\in\mathcal{J}$
@@
We begin by defining the equilibrium in the case where the CBAM value is fixed based on the initial vector of emission prices $\mathbf{t}$ and does not respond to subsequent price changes triggered by the introduction of CBAM, as discussed earlier in this Section. We then extend this definition to allow the CBAM value to adjust endogenously in response to the price shifts induced by the trade barrier.
@@
Our model conceptualizes the world economy as a network, where each node represents a sector in a country, and the links denote the flows of intermediate inputs between these sector-country pairs. One of the goals of our analysis is to evaluate the role of the network’s structure in mediating the effects of the CBAM. To ease the analysis of the model, we introduce additional notation.
To facilitate calibration, we define the steady state of the economy as the situation in which wedges $d_{ni}^{kj}$ and $\kappa_{ni}^{kj}$ are equal to zero for all combinations of $i$, $n$, $k$, $j$. Furthermore, we normalize the Hicks productivity parameter $A_{i}^{j}$ such that ${(1-\epsilon_{i}^{j})}^{\rho_{i}^{j}}{A_{i}^{j}}^{-(1-\rho_{i}^{j})}=1$ [^7] Note that this is equivalent to appropriately rewriting the normalization constant $\Upsilon_{i}^{j}$ in (5)..
@@
    $\alpha_{i}^{j}\equiv\frac{p_{i}^{j}c_{i}^{j}}{\sum_{j\in\mathcal{J}}p_{i}^{j}c_{i}^{j}}=\frac{p_{i}^{j}c_{i}^{j}}{I_{i}}=\frac{\chi_{i}^{j}(p_{i}^{j})^{1-\sigma}}{\sum_{j\in\mathcal{J}}\chi_{i}^{j}(p_{i}^{j})^{1-\sigma}}$
@@
5. Quantifying the effect of CBAM
In this section, we apply the model developed in the previous paragraphs to evaluate the effects of introducing the CBAM on trade across sectors and countries, emissions embodied in imports, carbon leakage, gross national expenditure, and real wages.
First, we solve the model in changes using the exact hat-algebra approach [dekle2008global], which enables us to evaluate different counterfactual scenarios. In the following exercises, we consider both adjustments: the full non-linear adjustments to a finite — though not necessarily marginal — policy shock, while maintaining computational efficiency through a reduced number of parameters. To illustrate the mechanisms at play, we provide a comparative static analysis of a marginal adjustment in CBAM in the Appendix A.2.
In our following exercises, we consider both an adjustment to the equilibrium with an exogenous CBAM and an endogenous CBAM, as explained in the previous section. That is, unlike standard fixed tariffs, the CBAM adjusts in response to domestic and foreign carbon prices – both of which may be influenced by the implementation of CBAM itself. This mechanism is intended to provide an incentive for faster convergence in clean technologies and regulations.
By considering both scenarios, we can assess the significance of the additional adjustment mechanism in both the case of a reduced CBAM, i.e., when the policy applies only to a subset of carbon-intensive goods, and the case of a full CBAM, i.e., when the policy applies to all sectors of the EU economy currently covered by the ETS. Throughout the remainder of the analysis, we assume that the reduced CBAM applies to the following six sectors in our data: Mining and Quarrying; Chemicals and Chemical Products; Non-Metallic Mineral Products; Basic Metals; Fabricated Metal Products; and Electricity, Gas, Steam, and Air Conditioning Supply.